Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-03-18
ISS On-Orbit Status 03/18/08

STS-123-1J/A Flight Day (FD) 8/9.

Crew sleep cycle today: Sleep 6:00am -2:30pm; wake 2:30pm -6:00am tomorrow.

EVA-3 was completed successfully by Rick Linnehan and Bob Behnken in 6h 53m, accomplishing most of its objectives.
(During the spacewalk, Linnehan (EV1) & Behnken (EV2) -

Installed the OTP (ORU {On-Orbit Replaceable Unit} Temporary Platform) and THA (Tool Holder Assembly) on the SPDM (Special Purpose Dexterous Manipulator),
Removed MLI (Multi-Layered Insulation) thermal blankets,
Installed the CLPA (Camera, Light & Pan/Tilt Assembly) on the SPDM,
Cleaned up & configured the SLP (Spacelab Pallet) for return (to be transferred with the SSRMS (Space Station Remote Manipulator System) from the POA (Payload ORU Attachment) to the Orbiter PLB (Payload Bay) tomorrow evening (FD9)),
Transferred the spare SSRMS yaw joint from the PLB to stowage on the ESP-2 (External Stowage Platform 2),
Transferred two spare DCSUs (Direct Current Switching Units) from the PLB to stowage on the ESP-2,
Transferred the LWAPA (Light Weight Adapter Plate Assembly) for installation on the Columbus EPF (External Payload Facility) and prepared for the installation of two MISSE-6 (Materials International Space Station Experiment) payloads,
Removed the MCAS EBCS (Mobil Common Attachment System External Berthing Camera System) cover flap as a get-ahead, and
Stowed the JLP (JEM Experiment Logistics Module Pressurized Section) trunnion covers on a handrail for future installation.
Tasks not completed:

MISSE-6B was not installed (its pip pins could not be properly aligned & locked for final installation on the LWAPA; MISSE-6B was returned to the PLB sidewall carrier),
MISSE-6A was not removed from the PLB carrier for installation due to lack of remaining EVA time (the MISSE-6 units will be added to EVA-5, probably replacing TBA-5 (Trundle Bearing Assembly 5) re-installation. EVA-4 will remain unchanged) , and
Some SPDM thermal MLI covers were left in place for later removal (viz., an EP blanket, the OTCM-2 (ORU Tool Changeout Mechanism 2) sock, and the flaps on SPDM joint 2-7).
Official start time of the spacewalk was 6:51pm EDT, about 32 min ahead of timeline, and it ended at 1:44am. Total EVA duration (PET = Phase Elapsed Time) was 6h 53min. It was the 107th spacewalk for ISS assembly & maintenance and the 79th from the station (57 from Quest, 22 from Pirs, 28 from Shuttle) totaling 486h 23min, the 11th for Expedition 16 (totaling 78h 31min) and the 8th so far this year. After today's EVA, a total of 135 spacewalkers (103 NASA astronauts, 21 Russians, and 11 astronauts representing Japan-1, Canada-4, France-1, Germany-2 and Sweden-3) have logged a total of 674h 45min outside the station on building, outfitting and servicing. It was also the 129th spacewalk involving U.S. astronauts.

After wakeup at ~2:30pm EST yesterday, ending the 8.5-hr sleep period before the spacewalk, the Airlock Crewlock (A/L CL) hatch was cracked at ~3:10pm for a hygiene break/with mask prebreathe for Linnehan and Behnken, after spending the night on 10.2 psi campout. Around 3:30pm, the hatch was closed again by IVs (Intravehicular Crewmembers) for EVA preparations in 10.2 psi, followed by EMU purge & prebreathe. Afterwards, IVs supported the CL depressurization and EV1/EV2 egress (EMUs switched to batteries at 6:51pm).

As part of pre-EVA activities, FE-1 Malenchenko powered down the ham radio equipment in SM (Service Module) and FGB at ~2:30am EST to prevent RF interference with the EMUs during the spacewalk.

Before her breakfast and EVA-support activities as IV, CDR Peggy Whitson completed another session with the SLEEP experiment (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) software for data logging and filling in questionnaire entries in the experiment's laptop session file on the HRF-1 laptop for downlink, as suggested on her discretionary 'job jar' task list. (To monitor the crewmember's sleep/wake patterns and light exposure, Peggy wears a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

Yuri Malenchenko made preparations for another run of the Russian/German TEKh-20 Plasma Crystal-3 Plus (PK-3+) experiment payload, the second of Expedition 16, by unstowing the hardware, installing it in the SM for operation and photographing the setup. The images were downlinked to TsUP via BSR-TM for inspection, and the FE-1 performed the initial leak check of the PK-3 Electronics Box before its evacuation. More work to come tomorrow. (The experiment is performed on plasma, i.e., fine particles charged and excited by HF (high frequency) radio power inside a vacuum work chamber. Main objective is to obtain a homogeneous plasma dust cloud at various pressures and particle quantities with or without superimposition of an LF (low frequency) harmonic electrical field. The experiment is conducted in automated mode. PK-3+ has more advanced hardware and software than the previously used Russian PKE-Nefedov payload.)

For Garrett Reisman's upcoming first session with the biomed experiment 'Integrated Immune', Peggy Whitson set up the equipment needed to support saliva collection. (Along with NUTRITION (Nutritional Status Assessment), Integrated Immune (Validating Procedures for Monitoring Crew member Immune Function) samples & analyzes participant's blood, urine, and saliva before, during and after flight for changes related to functions like bone metabolism, oxidative damage and immune function to develop and validate an immune monitoring strategy consistent with operational flight requirements and constraints. The strategy uses both long and short duration crewmembers as study subjects. The saliva is collected in two forms, dry and liquid. The dry samples are collected at intervals during the collection day using a specialized book that contains filter paper. The liquid saliva collections require that the crewmember soak a piece of cotton inside their mouth and place it in a salivette bag; there are four of the liquid collections during docked operations. The on-orbit blood samples are collected right before undocking and returned on the Shuttle so that analysis can occur with 48 hours of the sampling. This allows assays that quantify the function of different types of white blood cells and other active components of the immune system. Samples are secured in the MELFI (Minus-Eighty Laboratory Freezer for ISS). Also included are entries in a fluid/medications intact log, and a stress-test questionnaire to be filled out by the subject at begin and end. Urine is collected during a 24-hour period, conventionally divided into two twelve-hour phases: morning-evening and evening-morning.)

During the spacewalk, which MS2 Foreman supported as IV, Whitson and MS3 Doi continued configuring the JLP today. (Activities included rotation of the DMS (Data Management System) rack, retrieving two hard dummy panels from the rack standoff, assembling a dummy panel and installing it to the rack standoff. In preparation for Flight 1J, Takao & Peggy retrieved a JEM (Japanese Experiment Module) equipment bag as well as two JTVE (JEM Television Equipment) booms which for their assembly tomorrow (they will then remain in the JLP for their eventual deployment during Flight 1J's EVA-2.)

After the spacewalkers' ingress this morning at 1:44am, post-EVA activities by Whitson, Linnehan, Foreman and Behnken in the Airlock (A/L) consisted of recharging the EMU/spacesuits with water from PWR (Payload Water Reservoir), then reconnecting the LTAs (Lower Torso Assemblies) to the EMUs and capping the UIA (Umbilical Interface Assembly), initiating and monitoring the regeneration of the METOX (Metal Oxide) CO2 filter canisters, initiating the discharge/recharge process on the EMU batteries in the BSA (Battery Stowage Assembly), turning around the DCS-760 EVA camera (including downloading its photographs), and reconfiguring EVA tools.

After the EVA, Eyharts & Reisman maneuvered the SSRMS to re-grapple the SPDM and feed operational power to it for its heaters. The SPDM LEE (Latching End Effector) was then commanded to release the SLP, after which Dextre was maneuvered by SSRMS to an overnight park position.

Tonight at ~5:00pm, Eyharts & Reisman, joined by PLT Gregory Johnson, will command both SPDM arms in stow position and, at ~7:30pm, relocate Dextre for stowage on the U.S. Lab PDGF (Power & Data Grapple Fixture).

After releasing Dextre, Garrett and Greg will move the SSRMS over to grapple the SLP, detach it from the POA and transfer it to the Orbiter for berthing in the PLB. This activity should be completed by ~12:00midnight.

FE-1 Malenchenko hooked up the necessary 'plumbing' and compressor for transferring potable water from the BV1 tank of Progress M-63/28P to an EDV storage container. (The water transfer, monitored by Yuri several times during the 'day', at first was by self-flow (under its own tank pressure), then using the compressor pump via a GZhS gas/liquid separator, to remove air bubbles in the water. The BV2 Rodnik tank was emptied of its water by Yuri on 2/14. Filling of the empty Progress BV1 & BV2 tanks with urine will be scheduled later.)

Yuri conducted the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various Russian segment (RS) hatchways, including the FGB-to-Soyuz tunnel, and the FGB-to-Node passageway. (This is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).)

The FE-1 also completed the routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables.

FE-2 Eyharts performed periodic maintenance on the U.S. OGS (Oxygen Generation System), activating the WDS (Water Delivery System) from a filled PWR (Payload Water Reservoir), then deactivating it, leaving the PWR stowed in front of it. (Like the Russian Elektron, OGS produces O2 from water by electrolysis, dumping the also generated H2 (hydrogen) through venting.)

Malenchenko spent several hours on transferring discarded equipment and waste to Progress 28P and loading the spacecraft-turned-trash can preparatory to its separation on 4/1 for destructive reentry, tracking transfers in the IMS (Inventory Management System).

Reisman, as Eyharts before him, had 60 minutes for himself for general orientation (station familiarization & acclimatization), as is standard daily rule for the first two weeks after starting station residency.

Later, before sleeptime this morning (6:00am), Yuri conducted the daily IMS maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Working from the Russian 'available time' suggestions list, Yuri Malenchenko performed the regular daily checkup on the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at +20 degC, including a temperature check on its ART (automatic temperature recorder).

A second task list item for the FE-1 for FD7 was the daily monitoring, picture-taking and downloading for the BIO-5 Rasteniya-2 ("Plants-2") experiment which researches growth and development of plants (peas) under spaceflight conditions in the Lada-12 greenhouse from IBMP (Institute of Bio-Medical Problems {Russian: IMBP}).

Despite their busy timelines, all ISS crewmembers had time scheduled yesterday (as well as tonight after wakeup) for conducting their regular physical exercise.

The ISS crew began their sleeptime this morning at 6:00am, the Shuttle at 6:30am. For both crews, currently asleep, workday begins later today at 2:30pm EDT.

KOB-2 TCS Pump Update: After some more testing by TsUP-Moscow, RSC-Energia has determined that micro pump #1 (one of two) in the 4SPN1 replaceable pump panel of the SM TCS (SOTR, Thermal Control Systems)'s KOB-2 loop is failed. Its replacement will be planned pending launch of spare hardware (not currently onboard). At present, thermal loop KOB-2 is active with its alternate twin-pump package 4SPN2, with KOB-1 available as redundant backup. (Each of the two SOTR KOB loops has two redundant pump panels with two redundant pumps each. While in the early years of Mir and ISS the pumps were integral to the SPN panels, the current design allows them to be replaced without requiring an entire new SPN block.)

ATV Update (Flight Days 7, 8, 9): No phasing or rendezvous burns were performed during FDs 7, 8, & 9. PDE-4 (Propulsion Drive Electronics #4) was successfully re-integrated on 3/15 (FD7) at 6:06am EDT as part of the recovery from the FDIR actions during the MC1-2 burn. The switch to CPF-2 (Communication Processor Formatter #2) was completed successfully on 3/17 (FD9) at 4:30am. The next scheduled burns are TV1_1 and TV1_2 today (3/18) at 6:30:34pm & 7:16:45pm. They will be performed with the OCS (main engines) and in helium (He) pressurant configuration B, which exercises the redundant set of He pressure regulators and is the first time these regulators will be used for an OCS burn. They have been used previously with the ACS (attitude control thrusters) and have performed nominally, as also during the initial post-insertion pressurization.

ISS Crew Sleep Shift Planning: To synchronize the ISS crew's timeline with STS-123/1J/A docking and subsequent docked activities, Peggy's, Yuri's and Leo's wake/sleep cycle is undergoing a number of shifts which started on 3/11. For the next six days, the wake/sleep shift schedule is as follows (all times EDT):

FD9 Wake: 2:30pm (3/18) - 6:00am (3/19)
Sleep: 6:00am - 2:30pm (3/19)

Wake: 2:30pm (3/19) - 5:00am (3/20)
Sleep: 5:00am - 1:30pm (3/20)

Wake: 1:30pm (3/20) - 5:00am (3/21)
Sleep: 5:00am - 1:30pm (3/21)

Wake: 1:30pm (3/21) - 4:00am (3/22)
Sleep: 4:00am - 12:30pm (3/22)

Wake: 12:30pm (3/22) - 4:00am (3/23)
Sleep: 4:00am - 12:30pm (3/23)

Wake: 12:30pm (3/23) - 3:30am (3/24)
Sleep: 3:30am - 12:00pm (3/24)

No CEO photo targets uplinked for today.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use